# Discovery of the optical counterpart of the fast X-ray transient EP240414a

### v1

@@ 5 Summary and conclusions
By comparing the full sample of known-redshift EP-discovered FXTs with the historical XMM-Newton and Chandra samples of FXTs, we find that EP FXT events are consistently and systematically more luminous than all previously discovered FXTs. EP 240414a is the most nearby of the EP-discovered FXT sample to date by a considerable margin and although its redshift closely resembles the XMM-Newton and Chandra samples, it is significantly more luminous. EP is optimised for uncovering volumetrically rare, high-redshift, X-ray-luminous events, many of which appear to be linked to gamma-ray bursts; these FXTs appear to be distinct from the low redshift, historical sample of FXTs uncovered by XMM-Newton and Chandra.
@@
Acknowledgments
SJS, SS, KWS and DRY acknowledge funding from STFC Grants ST/Y001605/1, ST/X001253/1, ST/X006506/1 and ST/T000198/1. SJS acknowledges a Royal Society Research Professorship and the Hintze charitable foundation. HFS is supported by Schmidt Sciences, LLC. MN is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 948381) and by UK Space Agency Grant No. ST/Y000692/1. FC acknowledges support from the Royal Society through the Newton International Fellowship programme (NIF/R1/211296). AJC acknowledges support from the Hintze Family Charitable Foundation. TWC & AA acknowledge the Yushan Fellow Program by the Ministry of Education, Taiwan for the financial support (MOE-111-YSFMS-0008-001-P1). SY acknowledges the funding from the National Natural Science Foundation of China under Grant No. 12303046 and the Henan Province High-Level Talent International Training Program. A.B is grateful to the Royal Society, United Kingdom and also acknowledges SERB (SB/SRS/2022-23/124/PS) for financial support.

### v2

@@ 6 Summary and conclusions
By comparing the full sample of known-redshift EP-discovered FXTs with the historical XMM-Newton and Chandra samples of FXTs, we find that EP FXT events are consistently and systematically more luminous than all previously discovered FXTs. EP240414a is the most nearby of the EP-discovered FXT sample to date by a considerable margin and although its redshift closely resembles the XMM-Newton and Chandra samples, it is significantly more luminous. EP is optimised for uncovering volumetrically rare, high-redshift, X-ray-luminous events, many of which appear to be linked to gamma-ray bursts.
@@
Acknowledgments
The authors thank the anonymous reviewer for detailed comments that improved the quality of the manuscript.
SJS, SS, KWS and DRY acknowledge funding from STFC Grants ST/Y001605/1, ST/X001253/1, ST/X006506/1 and ST/T000198/1. SJS acknowledges a Royal Society Research Professorship and the Hintze charitable foundation. LR acknowledges support from the Canada Excellence Research Chair in Transient Astrophysics (CERC-2022-00009). HFS is supported by Schmidt Sciences, LLC. MN is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 948381) and by UK Space Agency Grant No. ST/Y000692/1. FC acknowledges support from the Royal Society through the Newton International Fellowship programme (NIF/R1/211296). AJC acknowledges support from the Hintze Family Charitable Foundation. TWC & AA acknowledge the Yushan Fellow Program by the Ministry of Education, Taiwan for the financial support (MOE-111-YSFMS-0008-001-P1). SY acknowledges the funding from the National Natural Science Foundation of China under Grant No. 12303046 and the Henan Province High-Level Talent International Training Program. A.B is grateful to the Royal Society, United Kingdom and also acknowledges SERB (SB/SRS/2022-23/124/PS) for financial support.
[…]
Appendix A Optical counterpart follow-up
A.1 Historical non-detections of an optical excess in ATLAS and Pan-STARRS
A.2 Optical photometric follow-up
A.3 Swift X-ray analysis
Appendix B Color evolution of AT 2024gsa
Here we include a visual demonstration of the intrinsically red color of AT 2024gsa, compared with AT 2018cow (Prentice et al., 2018), GRB 071010A (Covino et al., 2008) and SN 1998bw (Clocchiatti et al., 2011). It is apparent that AT 2024gsa is significantly redder than the LFBOT AT 2018cow and also SN 1998bw at later times. The early color evolution of AT 2024gsa is somewhat similar to SN 1998bw, however the luminosity is much higher relative to SN 1998bw at a similar epoch (see Section 4 for more details).
Figure 6: The ($r-i$) color evolution of AT 2024gsa at early times (phases A and B), campared to the ($g-r$) color evolution of AT 2018cow (Prentice et al., 2018), ($R-I$) color evolution of GRB 071010A (Covino et al., 2008), and ($V-R$) color evolution of SN 1998bw (Clocchiatti et al., 2011).